# Federated Graph Learning for Cross-Domain Recommendation

### v1

@@ 4.1 Experimental setup
We study the effectiveness of FedGCDR with 16 popular domains on a real-world dataset Amazon [49]. To study the impact of the number of domains on model performance, we divide these domains into three subsets containing 4, 8, and 16 domains respectively and denote them as Amazon-4, Amazon-8, and Amazon-16 respectively. We filter the original data in different ways, and more details are given in Appendix C.1. In our experiments, Books and CDs are selected as the target domains. For the ratings in each domain, we first convert them to implicit data, where entries corresponding to existing user-item interactions are marked as 1 and others are marked as 0.
@@
We compare FedGCDR with the following state-of-the-art models: (1) FedGNN [33] is an attempt to adopt FL graph learning to recommender systems. Its recommendation performance could represent the data quality of the target domain and reflect negative transfer. In Tables 1 and 2, in order to distinguish FedGNN from the CDR baselines, we denote it by Single Domain. (2) EMCDR [15] is a conventional embedding-mapping CDR framework. We adjust it to the HFL framework following [6]. (3) PriCDR [17] is a privacy-preserving CDR framework, which adopted DP on the rating matrices rating matrix to ensure privacy. (4) FedCT [6] is a VAE-based federated framework that is the first attempt to protect intra-domain privacy in cross-domain recommendations. (5) FedCDR [7] is a dual-target federated CDR framework, where the user embeddings are transferred as knowledge to enhance the other domain’s model training. To adapt to the BS-CDR scenarios, we modify FedCDR by applying embedding averaging when receiving source domain embeddings.
Table 1: The recommendation performance on Amazon@Books. The best result for the same setting is marked in bold and the second best is underlined. Single Domain represents FedGNN and its performance is exactly the same on three sub-datasets. FedGCDR-DP is a complete implementation of our method while FedGCDR does not incorporate Gaussian noise.
@@
We provide implement details in Appendix C.2. and additional experimental results in Appendix D.
4.2 Recommendation performance
We report the model performance results in Tables 1 and 2. Single domain shows that the Book domain has better single-domain recommendation accuracy than the Music domain, which represents higher data quality and quantity. Under BS-CDR settings, FedGCDR outperforms all CDR baselines on all three sub-datasets, which confirms the effectiveness of the proposed model on real-world data.
To further study our model capacity in alleviating negative transfer, we first define two types of negative transfer: (1) Soft Negative Transfer (SNT), where recommendation models’ performance under the multi-domain setting is worse than that under the single-domain setting. This means that the knowledge from source domains poisoning the target domain’s model training. (2) Hard Negative Transfer (HNT), where recommended performance of a large number of source domains is lower than that of a small number of source domains. This means that the newly added domains are not conducive to the training of the target domain or conflict with the already added source domain.
Taking the Books domain as the target domain, EMCDR, PriCDR, FedCT and FedCDR both have serious negative transfer problems and lower performance on the three data subsets. From the SNT perspective, their performances is much worse than that of Single Domain as shown in Figure 4. From the HNT perspective, their performances under 16-domain settings is worse than that under the 8-domain and 4-domain settings, which suggests it is not appropriate to recklessly transfer knowledge to a well-informed domain. Our FedGCDR model successfully alleviates NT with consistently best and stable performance results. For the CDs domain, which has lower-quality data, the performance of the CDR models greatly improves with less NT in Figure 4. From the SNT perspective, negative transitions now has a lighter impact for all models. From the HNT perspective, on Amazon-8, the performance of all models decrease, which we believe is due to the strong negative knowledge generated by the four additional domains on top of Amazon-4. On Amazon-16, all methods achieve best performance which indicates more knowledge from the source domains helpd improve the model performance in the CDs domain. Overall, the capability of EMCDR, PriCDR and FedCDR to alleviate negative transfer is much higher than that of FedCT. This is because the proportion of target domain features in the final feature is guaranteed by tuning hyper-parameters to control the transfer ratio of the source domain. Meanwhile, FedGCDR avoids this kind of human involvement and maintains performance optimality on three sub-datasets. In conclusion, our experiments show the superiority of FedCDR in recommendation performance and the effectiveness of alleviating NT.
@@ 4.3 Ablation study
To study the contribution of each module of FedGCDR, we implement two model variants, FedGCDR-M and FedGCDR-T. FedGCDR-T transfers the source domain embeddings without mapping. FedGCDR-M replaces the attention graph expansion with the average sum of source domain embeddings and removes the fine-tuning. We experiment with Books and CDs as target domains on the Amazon-16 dataset. The experimental results are shown in Figure 5. We make the following observations: (1) The two variants perform differently on different target domains. On the Books domain, FedGCDR-T performs better than FedGCDR-M, which indicates that for domains with higher data quality, preventing the transfer of negative knowledge from other domains is more important than mapping this knowledge better, and the Positive Knowledge Activation module meets the requirements of such domains. On the CDs domain, FedGCDR-M performs better than FedGCDR-T, which indicates that for domains that are deficient in information, mapping knowledge correctly is more important than preventing inter-domain negative knowledge, and the Positive Knowledge Transfer module meets these requirements. (2) Compared to FedGCDR, the absence of either module can cause a significant drop in performance. This indicates that in cross-domain recommendation, we should not only focus on transferring positive knowledge, but also control the spread of negative knowledge to the target domain, especially when a large number of domains.
@@ 5 Limitations
Our experiments were conducted on 16 domains of the Amazon dataset. While this extensive dataset covers broader source domains, relying on a single dataset may limit the generalizability of our model to data from other sources. Our approach uses overlapping users as a cross-domain bridge. Indeed, there are no widely-recognized cross-domain recommendation (CDR) datasets with more than three domains, aside from the Amazon dataset. Despite this limitation, we firmly believe that the significant improvements in privacy preservation and model performance demonstrated by FedGCDR underscore its superiority.
@@ 6 Conclusion
We proposed FedGCDR, a federated graph learning framework designed for BS-CDR. FedGCDR addresses the critical challenge of privacy preservation and negative transfer by employing a positive knowledge transfer module and a positive knowledge activation module. Our privacy-preserving method achieves best recommendation quality results on 16 domains of the Amazon dataset. In the future, we aim to extend FedGCDR to improve the recommendaiton performance of both the target and the source domains.
@@ Appendix A Privacy analysis
Due to the algorithmic nature of GNN, the source domain embeddings we pass are a function result on the user embeddings and item embeddings. This means that in the event of a successful inference attack, our user item interaction matrix is exposed to the threat of privacy disclosure. So we use DP to further protect privacy.
@@
$\Lambda=\frac{1}{1+\frac{1}{U}\sum Leak_{u}+\frac{1}{V}\sum Leak_{v}}.$ (16)
@@ Appendix B Cost analysis
Due to the complexity of the FedGCDR pipeline, we perform a theoretical analysis of the computational and communication cost of FedGCDR accordance with the HVH pipeline including horizontal source domain training, vertical positive knowledge tranfer module and horizontal positive knowledge activation module.
@@ B.1 Computational cost
Given a GAT model, let $V$ be the node set, $E$ be the edge set, and $F$ the embedding size. The computational cost of one propagation layer of classic GAT framework is $O(|V|FF^{\prime}+|E|F^{\prime})$ [37]. In the horizontal source domain training, our model is a simplified GAT variants which discard feature transformation and non-linear activation. For a $N^{K}$ layer model, the simplified computational cost is $O(N^{K}|E|F)$. In the vertical positive knowledge tranfer module, space mapping is carried by a $N^{m}$ layers’ MLP with computational cost $O(N^{m}F^{2})$. In the horizontal positive knowledge activation module, the first part is the simplified GAT model and the second part is the fine-tuning model with computational cost $O(F^{2})$. In conclusion, for the FedCDR framework with $N^{\mathcal{D}}$ domains, $T^{G}$ GAT-based federated model traing epochs, and $T^{F}$ fine-tuning epochs, the total computational cost is $O(T^{G}(N^{\mathcal{D}}N^{K}|E|F+N^{m}F^{2})+T^{F}F^{2}))$. Cause $N^{m}F^{2}\ll N^{\mathcal{D}}N^{K}|E|F$, we get the final computational cost $O(T^{G}N^{\mathcal{D}}N^{K}|E|F+T^{F}F^{2}))$.
@@ B.2 Communication cost
In FedGCDR, the global model and item embeddings are held by the domain server. Let $I$ be the item set and $F$ be the embedding size. The space complexity of global model and item embeddings are $O(F)$ and $O(|I|F)$ respectively. In the horizontal source domain training, the domain server distributes the global model and item embeddings and get the gradient with the same size. The communication cost is $O(F+|I|F)$. In the vertical positive knowledge tranfer module, the $N^{m}$ layers’ MLP and its gradients are transmitted with the commnication cost $O(N^{m}F^{2})$. In the horizontal positive knowledge activation module, the target domain additionally perform a fine-tuning stage with comunication cost $O(|I|F)$. In conclusion, for the FedCDR framework with $N^{u}$ users, $T^{G}$ GAT-based federated model traing epochs, and $T^{F}$ fine-tuning epochs, the total communication cost is $O(T^{G}N^{u}(N^{m}F^{2}+F+|I|F)+T^{F}N^{u}|I|F)$. Cause $N^{m}F^{2}+F\ll|I|F$, we get $O(N^{u}|I|F(T^{G}+T^{F}))$. According to the expression, the communication cost of our FedGCDR is basically equivalent to the cost of two HFL progress. The cost is reduced because knowledge transfer takes place in user space, thus avoiding large-scale information exchange.
@@ C.1 Dataset details
The Amazon dataset we used is the 2018 version and can be easily accessd in https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/. The basis of our domain selection strategy is the amount of data before performing data filtering. Thus, we sorted the domains contained in the Amazon dataset based on the amount of data in descending order and selected the top 16 domains. Similarly, Amazon-4 and Amazon-8 were selected accordingly. The only exception is that we prioritized the Movie domain, which has a relatively small amount of source data, based on popularity. In addition to multi-domain experiments, we randomly selected 2500 overlapping users in the Books domain and CDs domain to construct the dataset Amazon-Dual, so as to validate the performance of our FedGCDR in the conventional dual-domain scenarios where users full-overlap. The statistics of sub-datasets and processing details are shown in Table 3 and Table 4. The bottleneck time of FedGCDR is the federated-GAT training time in each domain, and we also show it in Table 4.
Dataset $|U|$ $|U_{d}|$ $|I_{d}|$ $|R_{d}|$ avg (sparsity)
@@ Appendix D Additional experimental results
D.2 Dual-domain scenario
Table 5: Dual-domain CDR performance.
Model Books $\rightarrow$ CDs Books $\leftarrow$ CDs
HR@10 NDCG@10 HR@10 NDCG@10
Single Domain 0.2713 0.1429 0.2594 0.1524
EMCDR 0.2816 0.1409 0.2596 0.1540
PriCDR 0.2903 0.1446 0.2662 0.1583
FedCT 0.2384 0.1239 0.2570 0.1551
FedCDR 0.2566 0.1376 0.2657 0.1554
FedGCDR-DP 0.3076 0.1552 0.2749 0.1602
FedGCDR 0.3323 0.1838 0.2958 0.1797
According to the experimental results shown in Table 5, our FedGCDR achieved the best experimental metrics in both knowledge transfer directions. This shows that our approach is also suitable for traditional scenarios where users full-overlap and have only a single source domain and a single target domain.
D.3 Privacy budget
@@
To study the effects of privacy budget $\epsilon$ on the model performance, we vary the privacy budget $\epsilon=\{4,8,16,32,64\}$ to affect the $\sigma$. We experimented on Amazon-16 with CDs as the target domain and fix $\delta=10{-5}$. We report the results Figure 7. From that we can observe that the model’s performance decreases as $\epsilon$ decreases, but the model performance is not completely destroyed by Gaussian noise. Thus, there is a trade-off between accuracy and privacy, where a smaller $\epsilon$ value adds more noise to embeddings for stronger privacy preservation but leads to more prediction error. Therefore, to balance the data privacy preservation capacity and the model performance, we set it as $\epsilon=8$.

### v2

@@ 4.1 Experimental setup
We study the effectiveness of FedGCDR with 16 popular domains on a real-world dataset Amazon [52]. To study the impact of the number of domains on model performance, we divide these domains into three subsets containing 4, 8, and 16 domains respectively and denote them as Amazon-4, Amazon-8, and Amazon-16 respectively. The statistics of sub-datasets are shown in Table 1. We filter the original data in different ways, and more details are given in Appendix C.1. In our experiments, Books and CDs are selected as the target domains. For the ratings in each domain, we first convert them to implicit data, where entries corresponding to existing user-item interactions are marked as 1 and others are marked as 0.
Dataset $|U|$ $|U_{d}|$ $|I_{d}|$ $|R_{d}|$ avg (sparsity)
@@
We compare FedGCDR with the following state-of-the-art models: (1) FedGNN [34] is an attempt to adopt FL graph learning to recommender systems. Its recommendation performance could represent the data quality of the target domain and reflect negative transfer. In Tables 2 and 3, in order to distinguish FedGNN from the CDR baselines, we denote it by Single Domain. (2) EMCDR [15] is a conventional embedding-mapping CDR framework. We adjust it to the HFL framework following [6]. (3) PriCDR [19] is a privacy-preserving CDR framework, which adopted DP on the rating matrices rating matrix to ensure privacy. (4) FedCT [6] is a VAE-based federated framework that is the first attempt to protect intra-domain privacy in cross-domain recommendations. (5) FedCDR [7] is a dual-target federated CDR framework, where the user embeddings are transferred as knowledge to enhance the other domain’s model training. To adapt to the BS-CDR scenarios, we modify FedCDR by applying embedding averaging when receiving source domain embeddings.
We provide implement details in Appendix C.2.
4.2 Recommendation performance
We report the model performance results in Tables 2 and 3. Single domain shows that the Book domain has better single-domain recommendation accuracy than the Music domain, which represents higher data quality and quantity. Under BS-CDR settings, FedGCDR outperforms all CDR baselines on all three sub-datasets, which confirms the effectiveness of the proposed model on real-world data.
To further study our model capacity in alleviating negative transfer, we first define two types of negative transfer: (1) Soft Negative Transfer (SNT), where recommender models’ performance under the multi-domain setting is worse than that under the single-domain setting. This means that the knowledge from source domains poisoning the target domain’s model training. (2) Hard Negative Transfer (HNT), where recommended performance of a large number of source domains is lower than that of a small number of source domains. This means that the newly added domains are not conducive to the training of the target domain or conflict with the already added source domain.
Taking the Books domain as the target domain, EMCDR, PriCDR, FedCT and FedCDR both have serious negative transfer problems and lower performance on the three data subsets. From the SNT perspective, their performances is much worse than that of Single Domain as shown in Figure 4. From the HNT perspective, their performances under 16-domain settings is worse than that under the 8-domain and 4-domain settings, which suggests it is not appropriate to recklessly transfer knowledge to a well-informed domain. Our FedGCDR model successfully alleviates NT with consistently best and stable performance results. For the CDs domain, the performance of the CDR models greatly improves with less NT in Figure 4.From the SNT perspective, information-poor domains have a lower probability of negative transfer, as they are inherently less well-trained. From the HNT perspective, on the Amazon-8 dataset, the performance of all models declines, which we attribute to the strong negative knowledge introduced by the four additional domains compared to the Amazon-4 dataset. On the Amazon-16 dataset, all methods achieve best performance which indicates more knowledge from the source domains can improve the model performance in the CDs domain. Overall, the capability of EMCDR, PriCDR and FedCDR to alleviate negative transfer is much higher than that of FedCT. This is because the proportion of target domain features in the final feature is guaranteed by tuning hyper-parameters to control the transfer ratio of the source domain. Meanwhile, FedGCDR avoids this kind of human involvement and maintains performance optimality on three sub-datasets. In conclusion, our experiments show the superiority of FedCDR in recommendation performance and the effectiveness of alleviating NT.
Table 2: The recommendation performance on Amazon@Books. Single Domain represents FedGNN and its performance is exactly the same on three sub-datasets. FedGCDR-DP is a complete implementation of our method while FedGCDR does not incorporate Gaussian noise. (The best result for the same setting is marked in bold and the second best is underlined. These notes are the same to others.)
@@ 4.3 Ablation study
To study the contribution of each module of FedGCDR, we implement two model variants, FedGCDR-M and FedGCDR-T. FedGCDR-T transfers the source domain embeddings without mapping. FedGCDR-M replaces the attention graph expansion with the average sum of source domain embeddings and omits the fine-tuning stage. We experiment with Books and CDs as target domains on the Amazon-16 dataset. The experimental results are shown in Figure 5. We make the following observations: (1) The two variants perform differently on different target domains. On the Books domain, FedGCDR-T performs better than FedGCDR-M, which indicates that for domains with higher data quality, preventing the transfer of negative knowledge from other domains is more important than mapping this knowledge better (in other words, the quality of external information holds greater significance than its quantity), and the Positive Knowledge Activation module meets the requirements of such domains. On the CDs domain, FedGCDR-M performs better than FedGCDR-T, which indicates that for domains that are deficient in information, mapping knowledge correctly is more important than preventing inter-domain negative knowledge (in other words, the quantity of external information holds greater significance than its quality), and the Positive Knowledge Transfer module meets these requirements. (2) Compared to FedGCDR, the absence of either module can cause a significant drop in performance. This indicates that in cross-domain recommendation, we should not only focus on transferring positive knowledge, but also control the spread of negative knowledge to the target domain, especially when a large number of domains.
4.4 Dual-domain scenario
Table 4: Dual-domain CDR performance.
Model Books $\rightarrow$ CDs Books $\leftarrow$ CDs
HR@10 NDCG@10 HR@10 NDCG@10
Single Domain 0.2713 0.1429 0.2594 0.1524
EMCDR 0.2816 0.1409 0.2596 0.1540
PriCDR 0.2903 0.1446 0.2662 0.1583
FedCT 0.2384 0.1239 0.2570 0.1551
FedCDR 0.2566 0.1376 0.2657 0.1554
FedGCDR-DP 0.3076 0.1552 0.2749 0.1602
FedGCDR 0.3323 0.1838 0.2958 0.1797
According to the experimental results shown in Table 5, our FedGCDR achieved the best experimental metrics in both knowledge transfer directions. This shows that our approach is also suitable for dual-domain scenarios where users full-overlap and have only a single source domain and a single target domain.
We provide experimental results on privacy budget in Appendix D.2.
@@ 5 Limitations
Our experiments were conducted on 16 domains of the Amazon dataset. While this extensive dataset covers broader source domains, relying on a single dataset may limit the generalizability of our model to data from other sources. Our approach uses overlapping users as a cross-domain bridge. Indeed, there are no widely-recognized cross-domain recommendation datasets with more than three domains, aside from the Amazon dataset. Despite this limitation, we firmly believe that the significant improvements in privacy preservation and model performance demonstrated by FedGCDR underscore its superiority.
@@ 6 Conclusion
We proposed FedGCDR, a federated graph learning framework designed for BS-CDR. FedGCDR addresses the critical challenge of privacy preservation and negative transfer by employing a positive knowledge transfer module and a positive knowledge activation module. Our privacy-preserving method achieves best recommendation quality results on 16 domains of the Amazon dataset. In the future, we aim to extend FedGCDR to improve the recommendation performance of both the target and the source domains.
7 Acknowledgment
The research was supported by Natural Science Foundation of China (62272403).
@@ Appendix A Privacy analysis
Due to the algorithmic nature of GNN, the source domain embeddings we pass are a function result on the user embeddings and item embeddings. This means that in the event of a successful inference attack, our user item interaction matrix is exposed to the threat of privacy disclosure. We apply differential privacy (DP) to further safeguard embeddings, following the approach of Cai et al.[53].
@@
$\Lambda=\frac{1}{1+\frac{1}{|U|}\sum Leak_{u}+\frac{1}{|V|}\sum Leak_{v}}.$ (16)
@@ Appendix B Cost analysis
Due to the complexity of the FedGCDR pipeline, we perform a theoretical analysis of the computational and communication cost of FedGCDR accordance with the HVH pipeline including horizontal source domain training, vertical positive knowledge transfer module and horizontal positive knowledge activation module.
@@ B.1 Computational cost
Given a GAT model, let $V$ be the number of nodes, $E$ be the number of edges, and $F$ the embedding size. The computational cost of one propagation layer of classic GAT framework is $O(VFF^{\prime}+EF^{\prime})$ [38]. In the horizontal source domain training, our model is a simplified GAT variants which discard feature transformation and non-linear activation. For a $N^{K}$ layer model, the simplified computational cost is $O(N^{K}EF)$. In the vertical positive knowledge transfer module, space mapping is carried by a $N^{m}$ layers’ MLP with computational cost $O(N^{m}F^{2})$. In the horizontal positive knowledge activation module, the first part is the simplified GAT model and the second part is the fine-tuning model with computational cost $O(F^{2})$. In conclusion, for the FedCDR framework with $N^{\mathcal{D}}$ domains, $T^{G}$ GAT-based federated model training epochs, and $T^{F}$ fine-tuning epochs, the total computational cost is $O(T^{G}(N^{\mathcal{D}}N^{K}EF+N^{m}F^{2})+T^{F}F^{2}))$. Cause $N^{m}F^{2}\ll N^{\mathcal{D}}N^{K}EF$, we get the final computational cost $O(T^{G}N^{\mathcal{D}}N^{K}EF+T^{F}F^{2}))$.
@@ B.2 Communication cost
In FedGCDR, the global model and item embeddings are held by the domain server. Let $I$ be the number of items and $F$ be the embedding size. The space complexity of global model and item embeddings are $O(F)$ and $O(IF)$ respectively. In the horizontal source domain training, the domain server distributes the global model and item embeddings and get the gradient with the same size. The communication cost is $O(F+IF)$. In the vertical positive knowledge transfer module, the $N^{m}$ layers’ MLP and its gradients are transmitted with the communication cost $O(N^{m}F^{2})$. In the horizontal positive knowledge activation module, the target domain additionally perform a fine-tuning stage with communication cost $O(IF)$. In conclusion, for the FedCDR framework with $N^{u}$ users, $T^{G}$ federated model training epochs, and $T^{F}$ fine-tuning epochs, the total communication cost is $O(T^{G}N^{u}(N^{m}F^{2}+F+IF)+T^{F}N^{u}IF)$. Cause $N^{m}F^{2}+F\ll IF$, we get $O(N^{u}IF(T^{G}+T^{F}))$. According to the expression, the communication cost of our FedGCDR is basically equivalent to the cost of two HFL progress. The cost is reduced because knowledge transfer totally takes place in user space, thus avoiding large-scale information exchange.
@@ C.1 Dataset details
The Amazon dataset we used is the 2018 version and can be easily accessed in https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/. The basis of our domain selection strategy is the amount of data before performing data filtering. Thus, we sorted the domains contained in the Amazon dataset based on the amount of data in descending order and selected the top 16 domains. Similarly, the Amazon-4 and Amazon-8 datasets were selected accordingly. The only exception is that we prioritized the Movie domain, which has a relatively small amount of source data, based on popularity. In addition to multi-domain experiments, we randomly selected 2500 overlapping users in the Books domain and CDs domain to construct the dataset Amazon-Dual, so as to validate the performance of our FedGCDR in the conventional dual-domain scenarios where users full-overlap. The processing details are shown in Table 5. The bottleneck time of FedGCDR is the federated-GAT training time in each domain, and we also show it in Table 5.
@@ Appendix D Additional experimental results
D.2 Privacy budget
@@
To study the effects of privacy budget $\epsilon$ on the model performance, we vary the privacy budget $\epsilon=\{4,8,16,32,64\}$ to affect the $\sigma$. We experimented on the Amazon-16 with CDs as the target domain and fix $\delta=10{-5}$. We report the results Figure 7. From that we can observe that the model’s performance decreases as $\epsilon$ decreases.The degradation in model performance suggests that our approach struggles to counteract the effects of high-intensity noise in a large number of domains, but the model performance is not completely destroyed by Gaussian noise. Thus, there is a trade-off between accuracy and privacy, where a smaller $\epsilon$ value adds more noise to embeddings for stronger privacy preservation but leads to more prediction error. Therefore, to balance the data privacy preservation capacity and the model performance, we set it as $\epsilon=8$.